### v1

@@ 2.2 Big Data and Machine Learning for Smart Grid
The research work [31] explores the integration of big data analytics (BDA) in smart grids and renewable energy power stations, focusing on predicting system stability. A framework is developed and applied to a dataset of 60,000 instances using various machine learning methods, achieving high accuracy in predicting stability. Three machine learning models -a decision tree, random forest classifier, and conventional neural network (Deep learning) were utilized. The first model is related to the classification model, where the aim was to identify if the smart grid system will be considered stable or not. A penalized linear regression has been employed by the fourth machine learning model to estimate the stability differential equation root. Notably, the penalized linear regression model attains 96% accuracy, while convolutional neural networks (CNN) achieve 87% accuracy for classification. The main contribution lies in demonstrating the effectiveness of BDA in enhancing decision-making and system stability in smart grid environments, paving the way for future research on larger datasets and broader applications.
@@
The study contributes novel methods for wind power forecasting, abnormal data handling, and generating consumption profiles, facilitating the integration of renewable energy into community grids. Deployed in a pilot REC in Belgium, the modules aim to align consumption with local generation, enhancing coordination and benefiting from advantageous pricing mechanisms. On the E-Cloud pilot project data, the ENSEMBLE model has demonstrated the greatest forecasting performance. Its output is calculated as the average of the results of four different machine learning models: random forests, gradient boosting decision trees, multilayer perceptrons, and bi-directional LSTMs. The training procedure has been modified to account for the automated detection of anomalous wind power data samples, which has significantly increased forecasting accuracy. a method that uses dynamic time warping to create community members’ representative power usage profiles.
@@
In Figure 11, the actual wind power generation (in black) and the wind power forecast (obtained with the ENSEMBLE model) as a function of time are illustrated for a random day in the test set.
2.4 Forecasting Solar Radiation using Ensemble Feature Selection and Machine Learning
Precise prediction of solar radiation is vital for the reliable operation of power systems that rely heavily on photovoltaic generation. This study [33] assesses the performance of machine learning (ML) methods in predicting solar radiation. The algorithms evaluated include Support Vector Regression (SVR), Extreme Gradient Boosting (XGBT), Categorical Boosting (CatBoost), and Voting-Average (VOA). A novel approach is introduced to choose significant variables and their previous observations, hence improving the accuracy of predicting. The ensemble feature selection approach combines Pearson’s coefficient, mutual information, random forest, and relief to choose the most related endogenous and exogenous inputs, together with their historical observation values.The suggested method for estimating solar radiation consists of five important phases, as illustrated in Figure 12. Tested on real data from Salvador, Brazil, VOA outperforms other algorithms in all prediction time horizons. Contributions include comparing ML algorithms, introducing CatBoost for solar forecasting, and proposing an ensemble feature selection method. Results show improved accuracy and highlight the importance of selecting appropriate input variables and past observations for solar radiation forecasting. For the summer and winter datasets, VOA had the faster training speed (183.93 s and 14.76 s, respectively). Being more sophisticated, VOA combines SVR, XGBT, and CatBoost. Further research is recommended to optimize parameter selection and explore applications in different locations and forecasting domains.
@@ 2.4 Forecasting Solar Radiation using Ensemble Feature Selection and Machine Learning
2.5 Optimal Hybrid AI Method (Time Delay Neural Network (TDNN)) For Smart Grid Energy Management
In order to increase the effectiveness of energy management on a smart grid like Renewable Energy Community, authors of this research [34] provide an optimum hybrid AI strategy in this work. In order to predict the future values of the energy characteristics in the community, this technique uses a Time Delay Neural Network. Afterwards, a stochastic model predictive control uses these predictions to optimize community operations using an appropriate battery energy storage system management approach. The Mean Absolute Error for solar generation, total energy consumption, and common services is 1.60 kW, 2.15 kW, and 0.30 kW, respectively, based on the outcomes of forecasts made using a public dataset with a 24-hour prediction horizon. In comparison to its rivals, the model predictive management fed with such forecasts yields the highest revenue. When compared to using the exact same system of management without taking use of forecasts from a forecasting approach, the total income is enhanced by 18.72%.
@@ 2.6 Techniques for Detecting Transient Stability Using CNN-LSTM
When a disturbance arises, electric power systems are more vulnerable to instability because they function closer to their limitations. Problems with transient stability are particularly common. [35], the research introduces a novel CNN-LSTM model (shown in Figure 13) to detect transient stability in electric power systems using historical data events. Addressing complexities in power system operation, the model considers noise, delay, loss in measurement data, line outage, and variable renewable energy (VRE) integration scenarios. Achieving over 99% accuracy in identifying stable and unstable conditions, the proposed model outperforms conventional CNN and LSTM approaches in terms of computation time, averaging 190.4 seconds. By training on synthetic data and considering various system parameters, this model demonstrates promise for enhancing transient stability detection in real-world power systems, albeit with challenges in implementation and optimization. In terms of computation time and performance, the CNN-LSTM approach outperforms the CNN and convLSTM approaches.
@@
3 More Application of AI in Renewable Energy Domain
The energy industry’s significant development in the use of AI, IoT, and smart grid infrastructure control has led to a study of cyber-attacks and a rapid rise in the number of solutions [36]. When it comes to load balancing renewable energy, automation is particularly crucial since it ensures grid stability by matching distributed generation to load [37]. The actions of rogue or rouge actors combined with inadequate system design, implementation, or configuration pose a threat to load balancing and the smart grid as a whole. Before a cybercatastrophy occurs, it is imperative to perform a thorough investigation about the viability and urgency of these possible dangers [37].
@@ 3.1 Implementing AI to predict renewable power
Researchers and the energy business have paid close attention to big data as well as artificial intelligence (AI) in recent years. This interest stems from two key factors: (i) growing processing power and (ii) the exponential global data expansion [38]. It is suggested to use Artificial Neural Networks (ANNs) for modeling and sizing a clean energy community (CEC) that uses energy storage devices, electric car charging stations, and a PV-wind hybrid system to fulfill building district energy demand by the authors of reserach [38]. While the second ANN is used to estimate the grid energy indication factor, the first ANN is utilized to anticipate the energy performance indicators, which include the fulfilled load fraction and the utilization factor of the energy generated. ANNs are trained on a vast amount of data in all weather scenarios, with adjustable power system configurations, and with a range of electrical demands. They use advanced models of each CEC component to anticipate the year CEC energy performance directly, without doing any system dynamic simulations. The impact of each input on each output was assessed using the Garson algorithm. Compared to those employed in ANN training, optimized ANNs have a very high prediction accuracy of CECs since they are made up of a single hidden layer with twenty neurons.
@@ 3.2 Preventing Cyberattacks in Renewable Energy Field
The clever renewable energy to cyberattacks, opening a new avenue for hackers to exploit the weaknesses in renewable energy [39]. This clearly impacts smart renewable energy’s operational and economical aspects [39]. A number of contradicting Cyber-Physical Power Systems (CPPSs) have seen moderate modifications in new kinds of power systems due to advancements in communication networks, monitoring and smart control systems, as well as the widespread usage of Internet-based structures [40]. Cyber components and power portions are typically connected in these architectures. CPPSs deal with recently developed problems like as security, vulnerability, resilience, stability, and dependability. Assessing, evaluating, and offering remedies to lessen or resolve these issues heavily rely on precise modeling techniques and investigating the interplay processes related to Smart Grids’ (SGs’) cyber-security. Many studies focused on creating reliable methods for identifying and/or defending against cyberattacks. These included studies on detection mechanisms that use federated learning techniques, perfect or imperfect false data attacks, false data-driven or data-injection attacks, and designing resilient communication networks [40].
@@
The research article [41] introduces a novel data-driven approach for detecting false data injection (FDI) attacks in energy forecasting systems, crucial for the reliability of power systems amidst increasing reliance on renewable energy and information technology. Utilizing cross-validation, least-squares, and z-score metrics, the proposed mechanism effectively identifies intrusions without relying on system models or parameters, enhancing system resilience. Through experiments with tree-based wind power forecasting models, the approach successfully detects and removes corrupted data, restoring forecasting accuracy and generalizability.
@@
The increasing availability of energy storage devices and the growing use of intermittent renewable energy supplies are causing significant changes in power systems [42]. These developments prompt smart-grid operators to imagine a day when microgrids would facilitate peer-to-peer energy trading, which will result in the creation of transactive energy systems. Due in major part to their remarkable durability, blockchains have attracted a lot of interest from academics and business for their possible use in decentralized TES. The research [42] investigates blockchain-based Transactive Energy Systems (TES) and introduces a novel class of attacks targeting gateways between prosumers and the system. The decentralized TES’s general architecture is depicted in Figure 14.The study formulates a comprehensive model of blockchain-based TES, analyzes various threat scenarios, and proposes attack mitigation strategies. Through experiments using GridLAB-D and a private Ethereum network, the research demonstrates that even simple attacks can manipulate market equilibria, but mitigation measures like detection and gateway switching effectively counteract these threats, ensuring the integrity and resilience of the system.
@@ 3.3 Optimizing consumption of energy employing artificial intelligence
The subject matter ”Cost Savings” via AI-powered energy efficiency in smart systems is one of the advantages. Lower operating expenses and electricity bills are directly correlated with decreased energy use. A variety of methods are included in AI technology, such as predictive modeling, data analytics, and machine learning. AI-powered systems may learn from data trends, make wise judgments, and optimize energy use in real time thanks to their capabilities. When it comes to optimizing energy usage in smart buildings, AI has several advantages. It makes it possible to gather and analyze enormous volumes of data from different sensors, meters, and Internet of things devices. One of the main causes of buildings’ energy consumption is their HVAC (heating, ventilation, and air conditioning) systems. By taking into account variables like occupancy, the state of the weather, and the need for thermal comfort, AI-based algorithms can improve HVAC operations.
@@
[43] This study proposes a reliable method to optimize energy consumption in buildings using EnergyPlus software and a robust artificial neural network (ANN) model trained with data from a research center in Iran. Genetic algorithms are used to optimize a building’s energy usage. Highly efficient features are evaluated in the building’s energy usage. By conducting sensitivity analysis, the most influential parameters on energy consumption are identified, with the number of occupants and wall U-value being the most significant. The proposed ANN model accurately predicts energy consumption, enabling optimization through a Genetic Algorithm-based approach, resulting in a substantial 35% reduction in energy usage. Overall, the study provides valuable insights into optimizing energy consumption in buildings, offering a promising methodology for similar applications.
4 Challenges of AI adoption in the smart energy industry
@@
The smart energy industry faces a variety of bottleneck obstacles to AI adoption, including poor or nonexistent data, problems with AI network parameters, difficulties with technical infrastructure, a shortage of skilled professionals, integration difficulties, risks, and legal and compliance concerns. For building energy systems, failure detection and diagnosis provide additional difficult issues [44]. Various studies acknowledge that inadequate information and data insecurity are two of the main problems that energy systems face [45]. The performance and dependability of the energy system are impacted by subpar sensors, controllers, and controlled devices for data estimation and energy system operation. The energy market presents additional difficulties for the power grid’s intricate correlations and strong coupling, as well as for the high data dimensionality and enormous complexity of large-scale simulation grid data [46]. For grid operations, integrating renewable energy sources like solar and wind with AI is equally challenging and complex [47].
@@
The adoption of artificial intelligence (AI) in the smart energy industry faces several challenges. These include a lack of key AI skills among decision-makers, hindering the understanding of AI’s potential benefits [24]. Additionally, there is a shortage of practical expertise in developing reliable AI applications with real-world benefits, leading to reluctance among power companies to adopt new strategies. Outdated power system infrastructure [24]further complicates modernization efforts, with utility companies struggling to manage vast amounts of dispersed and disorganized data. Economic pressures also impede the integration of innovative energy technologies due to high costs and resource requirements. Furthermore, decentralization and diversification of energy supply, coupled with emerging AI technologies, present complex challenges for energy production and distribution worldwide. Dependency on cellular technologies limits AI’s potential [24], especially in developing economies. Moreover, AI-based applications often act as black boxes for consumers, raising concerns about transparency and security, particularly in the face of increasing cyber threats in the digital energy infrastructure. Addressing these challenges is essential to unlock the full potential of AI in revolutionizing the smart energy industry.
@@ 5 Current Challenges and Future Research Directions
The integration of artificial intelligence (AI) in renewable energy research poses various challenges and opens avenues for future exploration. Firstly, the complexity and cost of AI algorithms used in optimization, design, and management of renewable energy systems need to be addressed. Simplifying and making these models cost-effective is crucial for widespread adoption and scalability.
Addressing real-time issues such as automatically controlling load demand for wind and solar power farms under extreme weather conditions, enhancing the efficiency of geothermal plants using IoT and AI, automating circuit breakers and relays under fault conditions, and preventing/detecting the impact of lightning on transmission lines remain critical challenges.
Regulators and policymakers must recognize AI’s systemic impact on reducing greenhouse gas emissions and ensure flexible policies to facilitate the rapid deployment of AI technologies in the energy sector. Incorporating ethical implications of AI in corporate social responsibility standards is essential for responsible AI deployment. Establishing participatory mechanisms to understand stakeholder expectations and limitations in AI application cases is imperative. Building AI operating systems that accurately identify energy efficiency and cost savings for consumers and businesses will be invaluable.
Furthermore, enhancing transparency and accessibility of system data, protecting AI-based models from security risks, and integrating AI with other technologies like robotics, sensors, and IoT for optimization are key areas for future research. With distributed energy resources and microgrids, artificial intelligence (AI) may be a key component of the grid’s integration of renewable energy sources and promote operational independence. AI may be introduced to developing power markets with the help of international financial institutions, enabling a sustainable energy transition.
@@
From revolutionizing medical diagnoses [48], [49], [50], [51], [52] to enhancing personalized learning in education, and optimizing transportation networks for efficiency, artificial intelligence has been instrumental across various sectors [53], [54]. Now, with recent advancements, the power sector is harnessing AI’s potential to improve grid management, forecast energy demands, and enhance overall operational efficiency. Researchers anticipate that major gains in energy efficiency and sustainability will result from the integration of artificial intelligence (AI) with renewable energy sources. Artificial intelligence (AI) has the potential to increase forecasting accuracy, stabilize the grid, optimize renewable energy systems, and provide more intelligent energy management options. Given AI’s capacity for massive data analysis and real-time decision-making, we see a future in which renewable energy sources will predominate in supplying the world’s energy needs while minimizing its negative environmental effects. A more robust and sustainable energy environment is anticipated as a result of the convergence of AI with renewable energy, opening the door to a cleaner and greener future.
@@
[30] 2021 Used deep learning (DL) method AB-Net for forecasting of renewable energy generation Solar Dataset [55], Wind Dataset [56] AB-Net approach provides MSE 0.0004, RMSE 0.0189.
[31] 2022 Used decision tree, random forest, and conventional neural network, linear regression for smart grid management Smart Grid data [57] Linear regression model attains 96% accuracy.
[32] 2021 Developed a day-ahead wind power forecasting algorithm using Machine Learning algoriths (random forests, gradient boosting decision trees, multilayer perceptrons, and bi-directional LSTMs) to minimize electricity bills Meteo Data (Wind speed, pressure, temperature) [58], [32] Ensemble model provided improved RMSE 2327 kW.
[33] 2023 Extreme Gradient Boosting (XGBT), Support Vector Regression (SVR), Categorical Boosting (CatBoost), and Voting-Average (VOA), which combines SVR, XGBT, and CatBoost, are the machine learning techniques that are employed for forecasting of solar radiation Brazilian National Institute of Meteorology website (INMET) [59] The training speed of VOA was greater than that of XGBT. Winter-VOA forecasting accuracy was 0.2417 MAE, whereas summer-VOA forecasting accuracy was 0.2877 MAE.
[34] 2022 Time Delay Neural Network for energy management Data of PV generation plant [60] Total income is enhancedby 18.72%
[35] 2022 CNN-LSTM model to detect transient stability in electric power systems synthetic dataset generated from the DigSILENT PowerFactory simulation, available at: [61] Discovery of transient stability exceeded 99%.
[38] 2022 Artificial Neural Networks (ANNs) for electric vehicle (EV) charging station and wind-photovoltaic battery HRES performance assessment The computation of the eight dimensionless input data requires the input climate data from the 48 locales and the load data from the three building districts under consideration [38] The R2 values for the SLF, UF, and GEIF forecasts are, in order, 0.952, 0.925, and 0.942.
[42] 2021 explored TES based on blockchain and presented a new category of cyberattacks Historical data on the prosumers’ previous bids Threats reduced by switching the gateway and using detection.
[43] 2020 Genetic algorithms are used to optimize a building’s energy usage. Highly efficient features are evaluated in the building’s energy usage. After being gathered, weather information was loaded into the Metronome program (”National and Local Weather Radar, Daily Forecast, Hurricane and information from The Weather Channel and weather.com,” n.d.) Garson’s method demonstrates the ANN model’s excellent accuracy of 99.9% when it comes to remodeling.

### v2

@@ 2.2 Big Data and Machine Learning for Smart Grid
With an emphasis on system stability prediction, the research study [31] investigates the implementation of big data analytics (BDA) between intelligent grids and energy from renewable sources power plants. A framework is developed and applied to a dataset of 60,000 instances using various machine learning methods, achieving high accuracy in predicting stability. Three ML models were used: deep learning with a typical neural network, random forest classifier, as well as decision tree. The goal of the first model, which is connected to classification model, was to determine whether or not the smart grid system will be regarded as stable. The fourth ML model estimates the equilibrium differential equation root using a penalized linear regression. Notably, convolutional neural networks (CNN) obtain 87% accuracy for classification, compared to penalized linear regression model’s 96% accuracy. The main contribution lies in demonstrating the effectiveness of BDA in enhancing decision-making and system stability in smart grid environments, paving the way for future research on larger datasets and broader applications.
@@
The research introduces innovative approaches for wind power prediction, anomaly detection in data, and the generation of consumption profiles, facilitating the seamless integration of renewable energy into community grids. Implemented initially in a pilot Renewable Energy Community (REC) in Belgium, these modules aim to synchronize energy consumption with local generation, thereby improving coordination and leveraging favorable pricing mechanisms. Based on data from the E-Cloud pilot project, the ENSEMBLE model has exhibited superior forecasting accuracy. It aggregates predictions from four distinct machine learning models: random forests, gradient boosting decision trees, multilayer perceptrons, and bi-directional LSTMs, providing robust forecasting outcomes. The training procedure has been modified to account for the automated detection of anomalous wind power data samples, which has significantly increased forecasting accuracy. a method that uses dynamic time warping to create community members’ representative power usage profiles.
@@
In Figure 11, the graph shows the actual wind power generation (in black) alongside the wind power forecast (generated using the ENSEMBLE model) over time, depicting data from a random day within the test set.
2.4 Using Machine Learning and Ensemble Feature Selection to Forecast Solar Radiation
Precise prediction of radiation from solar is vital for the reliable operation of power systems that rely heavily on photovoltaic generation. This study [33] assesses the performance of machine learning (ML) methods in predicting solar radiation. The algorithms evaluated include Support Vector Regression (in short SVR), Categorical Boosting (in short CatBoost), Voting-Average (in short VOA) , and Extreme Gradient Boosting (known as XGBT). A novel approach is introduced to choose significant variables and their previous observations, hence improving the accuracy of predicting. The most related internal and extrinsic inputs, together with their historical observation values, are selected using an ensemble feature selection technique that incorporates Pearson’s coefficient, mutual information, RF, and alleviation. The suggested method for estimating solar radiation consists of five important phases, as illustrated in Figure 12. Tested on real data from Salvador, Brazil, VOA outperforms other algorithms in all prediction time horizons. Contributions include comparing ML algorithms, introducing CatBoost for solar forecasting, and proposing an ensemble feature selection method. Results show improved accuracy and highlight the importance of selecting appropriate input variables and past observations for solar radiation forecasting. For the summer and winter datasets, VOA had the faster training speed (183.93 s and 14.76 s, respectively). Being more sophisticated, VOA combines SVR, XGBT, and CatBoost. Further research is recommended to optimize parameter selection and explore applications in different locations and forecasting domains.
@@ 2.4 Using Machine Learning and Ensemble Feature Selection to Forecast Solar Radiation
2.5 Time Delay Neural Network (known as TDNN): The Best Hybrid AI Approach for Smart Grid Management of Energy
In their study [34], the authors propose an optimal hybrid AI strategy to enhance energy management efficiency in smart grids like Renewable Energy Communities. This approach employs a Time Delay Neural Network to predict future energy characteristics within the community. Subsequently, a stochastic model predictive control utilizes these predictions to optimize community operations through effective management of battery energy storage systems. Results from forecasting using a publicly available dataset show Mean Absolute Errors of 1.60 kW for solar generation, 2.15 kW for total energy consumption, and 0.30 kW for common services over a 24-hour prediction horizon. Comparative analysis reveals that the predictive management model significantly increases revenue compared to competitors. Specifically, leveraging forecasts improves total income by 18.72% compared to employing the same management system without utilizing forecasting.
@@ 2.6 Techniques for Detecting Transient Stability Using CNN-LSTM
When a disturbance arises, electric power systems are more vulnerable to instability because they function closer to their limitations. Problems with transient stability are particularly common. [35], the research introduces a novel CNN-LSTM model (shown in Figure 13) to detect transient stability in electric power systems using historical data events. The model addresses challenges in power system operation, taking into account factors such as noise, delays, data measurement losses, line outages, and scenarios involving the integration of variable renewable energy (VRE). Achieving over 99% accuracy in identifying stable and unstable conditions, the proposed model outperforms conventional CNN and LSTM approaches in terms of computation time, averaging 190.4 seconds. By training on synthetic data and considering various system parameters, this model demonstrates promise for enhancing transient stability detection in real-world power systems, albeit with challenges in implementation and optimization. In terms of computation time and performance, the CNN-LSTM approach outperforms the CNN and convLSTM approaches.
@@
3 Additional Applications of AI in the Renewable Energy Sector
The energy industry’s significant development in the utilization of AI, Internet of Things, and smart grid infrastructure control has led to a study of cyber-attacks and a rapid rise in the number of solutions [36]. When it comes to load balancing renewable energy, automation is particularly crucial since it ensures grid stability by matching distributed generation to load [37]. The actions of rogue or rouge actors combined with inadequate system design, implementation, or configuration pose a threat to load balancing and the smart grid as a whole. Before a cybercatastrophy occurs, it is really imperative to perform thorough investigation about viability and urgency of these possible dangers [37].
@@ 3.1 Implementing AI to predict renewable power
Researchers and the energy business have paid close attention to big data as well as artificial intelligence (AI) in recent years. This interest stems from two key factors: (i) growing processing power and (ii) the exponential global data expansion [38].The authors of the study [38] recommend employing Artificial Neural Networks (ANNs) to model and size a clean energy community (CEC) incorporating energy storage devices, electric vehicle charging stations, and a PV-wind hybrid system to meet district energy demands. One ANN is dedicated to estimating the grid energy indication factor, while another predicts energy performance indicators such as the fulfilled load fraction and utilization factor of generated energy. ANNs are trained extensively across diverse weather conditions, variable power system configurations, and a range of electrical demands, using advanced models for each CEC component to directly forecast annual energy performance without dynamic system simulations. The Garson algorithm assesses the impact of each input on the outputs. Optimized ANNs, characterized by a single hidden layer with twenty neurons, demonstrate highly accurate predictions for CECs compared to those used during training.
@@ 3.2 Preventing Cyberattacks in Renewable Energy Field
The clever renewable energy to cyberattacks, opening a new avenue for hackers to exploit the weaknesses in renewable energy [39]. This clearly impacts smart renewable energy’s operational and economical aspects [39]. Recent advancements in communication networks, monitoring, smart control systems, and the widespread adoption of Internet-based architectures have led to significant transformations in Cyber-Physical Power Systems (CPPSs), despite encountering some conflicting perspectives [40]. Cyber components and power portions are typically connected in these architectures. CPPSs deal with recently developed problems like as security, vulnerability, resilience, stability, and dependability. Assessing, evaluating, and offering remedies to lessen or resolve these issues heavily rely on precise modeling techniques and investigating the interplay processes related to Smart Grids’ (SGs’) cyber-security. Many studies focused on creating reliable methods for identifying and/or defending against cyberattacks. These included studies on detection mechanisms that use federated learning techniques, perfect or imperfect false data attacks, false data-driven or data-injection attacks, and designing resilient communication networks [40].
@@
The research paper [41] presents an innovative data-driven method for detecting false data injection (FDI) attacks in energy forecasting systems. This approach is critical to ensuring the reliability of power systems, particularly as reliance on renewable energy and information technology continues to grow. Utilizing cross-validation, least-squares, and z-score metrics, the proposed mechanism effectively identifies intrusions without relying on system models or parameters, enhancing system resilience. Through experiments with tree-based wind power forecasting models, the approach successfully detects and removes corrupted data, restoring forecasting accuracy and generalizability.
@@
The rising prevalence of energy storage devices and the expanded utilization of intermittent renewable energy sources are driving substantial transformations in power systems [42]. These developments prompt smart-grid operators to imagine a day when microgrids would facilitate peer-to-peer energy trading, which will result in the creation of transactive energy systems. Due in major part to their remarkable durability, blockchains have attracted a lot of interest from academics and business for their possible use in decentralized TES. The research [42] explores blockchain-based Transactive Energy Systems (TES) and introduces a new type of attack aimed at gateways between prosumers and the system. Figure 14 illustrates the general architecture of the decentralized TES. The study develops a detailed model of blockchain-based TES, examines various threat scenarios, and proposes strategies for mitigating these attacks. Experimental results using GridLAB-D and a private Ethereum network show that even basic attacks can disrupt market equilibria. However, mitigation measures such as detection and gateway switching are effective in countering these threats, thereby maintaining the system’s integrity and resilience.
@@ 3.3 Optimizing consumption of energy employing artificial intelligence
The subject matter ”Cost Savings” via AI-powered energy efficiency in smart systems is one of the advantages. Lower operating expenses and electricity bills are directly correlated with decreased energy use. A variety of methods are included in AI technology, such as predictive modeling, data analytics, and machine learning. AI-powered systems may learn from data trends, make wise judgments, and optimize energy use in real time thanks to their capabilities. When it comes to optimizing energy usage in smart buildings, AI has several advantages. It makes it possible to gather and analyze enormous volumes of data from different sensors, meters, and Internet of things devices. HVAC (heating, ventilation, and air conditioning) systems are a primary contributor to the energy consumption of buildings. AI-based algorithms can optimize HVAC operations by considering factors such as occupancy, weather conditions, and thermal comfort requirements.
@@
In this paper [43], a robust ANN network trained with data from an Iranian research center is combined with EnergyPlus software to suggest a dependable technique for optimizing energy usage in buildings. Genetic algorithms are used to optimize a building’s energy usage. Highly efficient features are evaluated in the building’s energy usage. Through the use of sensitivity analysis, the elements that have the greatest impact on energy consumption are determined; the most important ones are the wall U-value and the number of occupants. The proposed ANN model accurately predicts energy consumption, enabling optimization through a Genetic Algorithm-based approach, resulting in a substantial 35% reduction in energy usage. Overall, the study provides valuable insights into optimizing energy consumption in buildings, offering a promising methodology for similar applications.
4 Obstacles to Implementing AI in the Smart Energy Sector
@@
The smart energy sector encounters numerous obstacles in adopting AI, such as insufficient or poor-quality data, challenges with AI network parameters, technical infrastructure issues, a lack of skilled professionals, integration difficulties, as well as risks and legal compliance concerns. Additionally, failure detection and diagnosis in energy systems pose further complex problems [44]. Several studies highlight that insufficient information and data security are significant challenges for energy systems [45]. The performance and reliability of energy systems are adversely affected by inadequate sensors, controllers, and controlled devices used for data estimation and system operation. The energy market also faces complications due to the intricate relationships and strong interdependencies within the power grid, as well as the high dimensionality and complexity of large-scale simulation grid data [46]. Furthermore, integrating renewable energy sources such as solar and wind with AI in grid operations presents additional complexities [47].
@@
The integration of artificial intelligence (AI) in the smart energy sector encounters numerous obstacles. One significant issue is the lack of essential AI expertise among decision-makers, which impedes their comprehension of AI’s potential advantages [24]. Moreover, there is an evident gap in practical skills necessary for creating dependable AI applications that deliver tangible real-world benefits, causing power companies to hesitate in embracing new approaches. Efforts to modernize are also hampered by outdated power system infrastructure [24], with utility companies facing difficulties in managing extensive amounts of dispersed and unstructured data. Economic pressures also impede the integration of innovative energy technologies due to high costs and resource requirements. Furthermore, decentralization and diversification of energy supply, coupled with emerging AI technologies, present complex challenges for energy production and distribution worldwide. Dependency on cellular technologies limits AI’s potential [24], especially in developing economies. Furthermore, AI-based applications frequently function as opaque systems for consumers, leading to worries about transparency and security, especially given the rising cyber threats within the digital energy infrastructure. Overcoming these challenges is crucial to fully harness AI’s potential in transforming the smart energy sector.
@@ 5 Current Challenges and Future Research Directions
The incorporation of AI into research on renewable energy presents a number of obstacles as well as opportunities for further investigation. First and foremost, attention needs to be paid to the complexity and expense of AI algorithms employed in the design, optimization, and administration of renewable energy systems. Simplifying and making these models cost-effective is crucial for widespread adoption and scalability.
Critical challenges in real-time include automating circuit breakers and relays under fault conditions, detecting and preventing the impact of lightning on transmission lines, improving the efficiency of geothermal plants using IoT and AI, and regulating demand for load for wind and solar power farms under extreme weather.
Regulators and policymakers must recognize AI’s systemic impact on reducing greenhouse gas emissions and ensure flexible policies to facilitate the rapid deployment of AI in the domain of energy. Incorporating ethical implications of AI in corporate social responsibility standards is essential for responsible AI deployment. Establishing participatory mechanisms to understand stakeholder expectations and limitations in AI application cases is imperative. It will be extremely beneficial to develop AI applications that precisely detect energy cost and effectiveness reductions for both individuals and enterprises.
Future research should focus on improving system data accessibility and openness, safeguarding models based on artificial from security threats, and optimizing AI’s integration with additional technologies like robots, sensors, and the Internet of things. AI may play a significant role in the grid’s integrating sources of renewable energy and promotion of operational independence with distributed sources of energy and microgrids. AI may be introduced to developing power markets with the help of international financial institutions, enabling a sustainable energy transition.
@@
From revolutionizing medical diagnoses [48], [49], [50], [51], [52] to enhancing personalized learning in education, and optimizing transportation networks for efficiency, artificial intelligence has been instrumental across various sectors [53], [54]. Now, with recent advancements, the power sector is harnessing AI’s potential to improve grid management, forecast energy demands, and enhance overall operational efficiency. Researchers believe that combining AI with renewable energy sources will lead to significant improvements in sustainability and energy efficiency. AI has promise for improving forecasting precision, grid stabilization, renewable energy system optimization, and the provision of more intelligent managing energy choices. Given AI’s capacity for massive data analysis and real-time decision-making, we see a future in which renewable energy sources will predominate in supplying the world’s energy needs while minimizing its negative environmental effects. The combination of AI with renewable energy is expected to provide a more resilient and sustainable energy environment, paving the way for a more environmentally friendly future.
@@
khan et al. 2021 Used DL method AB-Net for forecasting of renewable energy generation Solar [55], Wind [56] MSE 0.0004, RMSE 0.0189.
mostofa et al. 2022 Used DT, RF, CNN, and linear regression for smart grid management Smart Grid [57] Linear regression 96% accuracy.
greve et al. 2021 Day-ahead wind power prediction using ML techniques (RF, GB, MLP, BiLSTMs) Meteo Data [58], [32] RMSE 2327 kW.
solano et al. 2023 XGBT, SVR, VOA, and CatBoost for solar radiation forecasting Brazilian INMET [59] Winter-VOA MAE 0.2417, Summer-VOA MAE 0.2877.
conte et al. 2022 Time Delay Neural Network for energy management PV generation plant data [60] Income enhanced by 18.72%.
azhar et al. 2022 CNN-LSTM model to detect transient stability in power systems DigSILENT synthetic dataset [61] Stability detection ¿99%.
mazzeo et al. 2022 ANN for EV charging station and HRES performance assessment Climate data from 48 locales [38] R2 values: SLF 0.952, GEIF 0.925, UF 0.942.
barreto et al. 2021 TES based on blockchain, exploring cyberattacks Prosumers’ historical bid data Threats reduced via gateway switching.
ilbeigi et al. 2020 Genetic algorithms to optimize building energy usage Weather data loaded into Metronome ANN model 99.9% accuracy in energy use remodeling.